OpenStudy (anonymous):

I don't understand how to do this at all.. Please help? y=6-x and y=x-2 a (2,4) b (-4,2) c (4,2) d no solution

OpenStudy (anonymous):

lets find y first: y=6-x we need to know x to get y, which we can get from our second equation y=x-2 solve y=x-2 for x

jimthompson5910 (jim_thompson5910):

y=6-x ... start with the first equation x-2=6-x ... replace y with x-2 (since y = x-2) now solve for x

OpenStudy (anonymous):

I'm so confused

jimthompson5910 (jim_thompson5910):

Do you see how I got to x-2=6-x ?

OpenStudy (anonymous):

yea...

jimthompson5910 (jim_thompson5910):

the goal now is to solve for x (ie isolate x)

jimthompson5910 (jim_thompson5910):

get x all by itself on one side

OpenStudy (anonymous):

Okay...

OpenStudy (anonymous):

We have two equations and two missing variables, x and y. y=6-x y=x-2 as you can see, y equals both 6-x, and x-2. That means we can do this: 6-x=x-2 now we only have 1 variable to solve for! This means that we can find out the actual value of x. do you know how to solve 6-x=x-2 for x?

OpenStudy (anonymous):

I have no clue on where to start... haha I'm sorry.. I've been trying to figure this out all day... :/

OpenStudy (anonymous):